### v2

# Revisiting Invariant Learning for Out-of-Domain Generalization on Multi-Site Mammogram Datasets

Abstract
Achieving health equity in Artificial Intelligence (AI) requires diagnostic models that maintain reliability across diverse populations. However, breast cancer screening systems frequently suffer from domain overfitting, degrading significantly when deployed to varying demographics. While Invariant Learning algorithms aim to mitigate this by suppressing site-specific correlations, their efficacy in medical imaging remains underexplored. This study comprehensively evaluates domain generalization techniques for mammography.
We constructed a multi-source training environment aggregating datasets from the United States (CBIS-DDSM, EMBED), Portugal (INbreast, BCDR), and Cyprus (BMCD). To assess global generalizability, we evaluated performance on unseen cohorts from Egypt (CDD-CESM) and Sweden (CSAW-CC). We benchmarked Invariant Risk Minimization (IRM) and Variance Risk Extrapolation (VREx) against a rigorously optimized Empirical Risk Minimization (ERM) baseline. Contrary to expectations, standard ERM consistently outperformed specialized invariant mechanisms on out-of-domain testing. While VREx showed potential in stabilizing attention maps, invariant objectives proved unstable and prone to underfitting. We conclude that engineering equitable AI is currently best served by maximizing multi-national data diversity rather than relying on complex algorithmic invariance.
I Introduction
Achieving health equity in Artificial Intelligence (AI) requires diagnostic models that maintain reliability across diverse patient populations and clinical environments. While deep learning has achieved expert-level performance in breast cancer screening [23, 13], these systems frequently suffer from domain overfitting—learning statistical shortcuts specific to a single hospital or demographic group. Consequently, performance often degrades significantly when models are deployed to under-represented populations or differing acquisition protocols, exacerbating existing healthcare disparities.
Recent empirical studies have highlighted the severity of this ”generalization gap.” For instance, Barros et al. [2] observed sharp performance declines when models trained on US cohorts were evaluated on Israeli populations, attributing the failure to variations in health policies and demographics. Similarly, De et al. [3] demonstrated that even commercial AI tools are fragile to changes in mammography equipment, often requiring local calibration to function correctly. These findings underscore a critical limitation: standard training methods (Empirical Risk Minimization) often latch onto ”spurious correlations” (e.g., scanner artifacts or site-specific tissue density distributions) rather than learning the causal biology of cancer.
To mitigate these disparities, Invariant Learning algorithms, such as Invariant Risk Minimization (IRM) [1] and Variance Risk Extrapolation (VREx) [9], have been proposed. Theoretically, these methods encourage models to ignore site-specific noise and focus on features that are stable across all environments. However, while applied to limited settings like lesion crops [21], the practical efficacy of these invariant penalties in high-dimensional, whole-mammogram analysis remains underexplored.
In this paper, we conduct a comprehensive evaluation of domain generalization techniques to engineer more equitable screening tools. Unlike previous studies limited to single-site or two-continent comparisons, we construct a diverse multi-source training environment aggregating public datasets from the USA, Portugal, and Cyprus. We benchmark invariant strategies against a rigorously optimized standard baseline on unseen out-of-domain cohorts from Egypt and Sweden. Our goal is to determine whether complex algorithmic invariance is necessary for robustness, or if a data-centric approach—maximizing the diversity of training data—remains the most effective path toward generalized breast cancer prediction.
II Methods
II-A Mammograms classification with tiled patches classifier initialization using CNNs
In the domain of whole mammogram classification, we build upon prior research that has demonstrated the effectiveness of CNNs [18, 16, 17]. In line with the approach presented in [18], our methodology involves initial training of the network using tiled patches, encompassing both lesion patches and background crops. Subsequently, the pre-trained network serves as the starting point for initializing the whole mammogram network. In this study, we employ a standard residual network (ResNet) [6], and a more modernized ConvNet (ConvNeXt) [11], which has demonstrated competitive accuracy and scalability compared to the Vision Transformer [4].
TABLE I: Comparative evaluation of Standard ERM against domain-invariant strategies (IRM, VREx) for whole-mammogram cancer classification. Experiments utilize ResNet34 and ConvNeXt-tiny backbones on both in-domain (CBIS-DDSM, EMBED) and out-of-domain (CDD-CESM, CSAW-$C$) test sets. Results indicate that Standard ERM remains highly competitive, particularly in OOD scenarios. (*: ERM approximated by setting the IRM penalty term to zero, which is usually used as a baseline in invariant learning papers).
Model Method In-Domain Out-of-Domain
CBIS-DDSM EMBED CDD-CESM CSAW-$C$
Acc AP AUC Acc AP AUC Acc AP AUC Acc AP AUC
ResNet34 ERM [20] 0.648 0.729 0.785 0.703 0.437 0.524 0.733 0.651 0.740 0.984 0.081 0.664
ERM* 0.566 0.745 0.788 0.738 0.559 0.630 0.692 0.508 0.652 0.979 0.032 0.549
IRM [1] 0.602 0.388 0.460 0.609 0.266 0.391 0.673 0.289 0.436 0.990 0.010 0.495
VREx [9] 0.570 0.741 0.797 0.641 0.521 0.641 0.644 0.545 0.672 0.914 0.059 0.628
ConvNeXt -tiny ERM [20] 0.733 0.742 0.797 0.814 0.719 0.746 0.794 0.731 0.832 0.979 0.108 0.727
ERM* 0.712 0.719 0.759 0.795 0.676 0.705 0.759 0.697 0.798 0.980 0.070 0.669
IRM [1] 0.401 0.527 0.590 0.326 0.513 0.604 0.618 0.577 0.709 0.989 0.019 0.580
VREx [9] 0.638 0.731 0.775 0.762 0.671 0.748 0.729 0.679 0.785 0.953 0.064 0.596
II-B Invariant Learning algorithms
II-B1 Invariant Risk Minimization (IRM)
Invariant Risk Minimization (IRM) [1] is a learning approach that focuses on estimating consistent correlations across diverse training distributions. By learning a data representation that ensures an optimal classifier performs well across all training distributions, IRM enables effective generalization to out-of-distribution scenarios. This capability is particularly valuable in real-world applications such as medical diagnosis, where the data distribution can vary significantly across different hospitals and institutions. The invariances captured by IRM are associated with the underlying causal structures that govern the data, enhancing its ability to uncover meaningful patterns and relationships.
Let $e\in\mathcal{E}$ be an environment with its own data distribution $\mathcal{X},\mathcal{Y}$. Denote $\mathcal{E}_{\mathrm{tr}}\subset\mathcal{E}$, $\mathcal{E}_{\mathrm{tr}}$ is an empirical set of environments or a set of environments that can be accessed. Let $\Phi:\mathcal{X}\rightarrow\mathcal{H}$ be a projection function of data into the latent space. $R^{e}(\Phi)$ will be a risk of environment $e$ under $\Phi$. The practical IRM objective is:
$\Phi^{*}=\underset{\Phi:\mathcal{X}\rightarrow\mathbb{R}^{d}}{\arg\min}\sum_{e\in\mathcal{E}_{\mathrm{tr}}}R^{e}(\Phi)+\lambda\left(\left.\nabla_{w}R^{e}(w\cdot\Phi)\right|_{w=1}\right)^{2}$ (1)
Here, $w$ is a simplified version of the predictor $g:\mathcal{H}\rightarrow\mathcal{Y}$ which is assumed to be optimal over all environments. To streamline the learning process, the value of $w$ is fixed at 1, allowing the network to concentrate on learning the representations captured by $\Phi$.
II-B2 Variance Risk Extrapolation (VREx)
Risk Extrapolation (REx) [9] demonstrates superior performance compared to Invariant Risk Minimization (IRM) when faced with scenarios where both causally induced distributional shifts and covariate shifts occur simultaneously. REx achieves this by effectively balancing its resilience to these types of shifts. In contrast, IRM assumes that the learned invariances from the training data alone are adequate for generalization to new domains, which may not always hold true. While both REx and IRM are capable of performing causal identification, REx demonstrates greater robustness in handling covariate shifts (changes in the input distribution). Covariate shift poses challenges when models are misspecified, when training data is limited, or when it fails to cover relevant areas of the test distribution.
REx is a robust optimization technique that operates within a perturbation set of extrapolated domains. Another variant, named VREx, introduces a penalty term that targets the variance of training risks. It has been shown that variants of REx can recover the causal mechanisms of the targets, while also providing some robustness to changes in the input distribution. VREx shares a similar objective to IRM but replaces the penalty term with the variance of training risks across different environments:
$\Phi^{*}=\underset{\Phi:\mathcal{X}\rightarrow\mathbb{R}^{d}}{\arg\min}[\lambda\left.Var(\{R^{e}(\Phi)\}_{e\in\mathcal{E}_{\mathrm{tr}}}\right)+\sum_{e\in\mathcal{E}_{\mathrm{tr}}}R^{e}(\Phi)]$ (2)
II-B3 Invariant Learning on multi-site datasets
In our study, we adopt the perspective that each dataset obtained from a particular hospital or institution represents a distinct environment. The training process involves straightforward computation of risks for each dataset, followed by the calculation of either the IRM or VREx objective. To ensure the validity of our approach, we ensure that the datasets used for training and testing originate from different sites or institutions. This approach enables us to evaluate the performance of our models across diverse environments and establish their robustness in handling variations between different data sources [5].
III Experiments
III-A Datasets
To ensure reproducibility and facilitate future benchmarking, this study exclusively utilizes publicly available datasets. We aggregated data from seven diverse sources: CBIS-DDSM (USA) [10], EMBED (USA) [7], INbreast (Portugal) [14], BMCD (Republic of Cyprus) [12], BCDR (Portugal) [15], CDD-CESM (Egypt) [8], and CSAW-CC (Sweden) [19].
Experimental Protocol: We designed our evaluation to assess both in-domain and out-of-domain (OOD) performance.
  • Training Domain: The primary training corpus consists of CBIS-DDSM, EMBED, INbreast, BMCD, and BCDR.
  • Out-of-Domain Testing: To evaluate generalization capability, we reserved the CDD-CESM and CSAW-CC datasets entirely for testing, as they originate from geographic regions not represented in the training phase.
  • In-Domain Testing: To monitor standard performance, we utilized the official test set of CBIS-DDSM and a held-out 25% subset of the EMBED dataset.
Strict patient-level separation was enforced across all splits to prevent data leakage.
III-B Experimental Settings
Architectures: We benchmark two Convolutional Neural Network (CNN) backbones: ResNet34 and ConvNeXt-tiny. While ResNet34 serves as a standard baseline, ConvNeXt-tiny was selected for its ability to achieve competitive accuracy and scalability comparable to Vision Transformers (ViTs) while maintaining the inductive biases of CNNs [11].
Training Strategy: To address the high resolution and data scarcity inherent in mammography, we adopt the two-stage transfer learning strategy proposed by Shen et al. [18].
  1. 1.
    Patch Pre-training: We first train a classifier on ROI-annotated patches (lesions and background) extracted from CBIS-DDSM [10], EMBED [7], INbreast [14], and BCDR [15]. This initializes the network with discriminative local features.
  2. 2.
    Whole-Mammogram Fine-tuning: The pre-trained weights are then transferred to the whole-mammogram classifier, which is fine-tuned on the full images.
Implementation Details: All models are optimized using Adam with a learning rate of $1\mathrm{e}{-3}$. Due to computational constraints, input mammograms are resized to $576\times 448$. We acknowledge that while higher resolutions (e.g., $1152\times 896$ [18] or larger [22]) often yield performance gains, our resolution choice balances resource efficiency with representational capacity. The quantitative results of these experiments are summarized in Table I.
III-C Quantitative Analysis
Table I details the performance of ResNet34 and ConvNeXt-tiny across in-domain and out-of-domain (OOD) settings. We compared four training paradigms: Standard ERM [20] (fully optimized), Approximated ERM* (IRM with $\lambda=0$), IRM [1], and VREx [9].
Robustness of Standard ERM: A critical finding of our study is the resilience of Standard ERM when trained on diverse data. On the unseen out-of-domain datasets (CDD-CESM and CSAW-CC), Standard ERM consistently achieved the highest generalization performance. With the ConvNeXt-tiny architecture, Standard ERM reached an AUROC of 0.832 on CDD-CESM and 0.727 on CSAW-CC, outperforming both the approximated baseline (ERM*) and the invariant methods. This highlights the importance of using a fully optimized Standard ERM as a rigorous benchmark when evaluating the added value of domain generalization techniques.
Instability of Invariant Constraints: Contrary to theoretical expectations, IRM consistently yielded the lowest performance, often trailing significantly behind the baselines. An analysis of training dynamics suggests severe underfitting; the complex penalty landscape of IRM appears difficult to traverse in high-dimensional mammography, leading to optimization instability. VREx showed better stability than IRM, occasionally matching ERM on in-domain tasks (e.g., ResNet34 on CBIS-DDSM), but it failed to provide a generalization benefit on the challenging OOD cohorts.
Architecture Impact: ConvNeXt-tiny consistently outperformed ResNet34 across all metrics and datasets. This performance gap was particularly pronounced on the EMBED test set, confirming that modernizing the CNN backbone provides a more tangible reliability improvement than altering the loss function.
III-D Qualitative Analysis
To interpret the decision-making process behind the quantitative results, we visualized Class Activation Maps (CAM) for the top-performing methods, ERM and VREx. Figures 1 through 2 illustrate discriminative regions identified by the models within the full mammograms.
Localization of Clinical Features: A key concern in medical AI is whether high performance stems from genuine pathology detection or spurious background correlations. Our visualizations suggest that Standard ERM achieves its superior quantitative performance by predominantly prioritizing clinical cues. As shown in the top rows of the Figure 1 and Figure 2, the ERM-trained models consistently focus attention on the specific lesion regions within the breast tissue. This demonstrates that despite lacking explicit invariance constraints, ERM is capable of learning robust, medically relevant features when trained on diverse multi-source data.
Attention Drift in Both Methods: While both ERM and VREx generally identify the correct region of interest in successful cases, neither method is immune to attention drift. We observe instances in both training paradigms where the model focuses on irrelevant areas, such as healthy fibroglandular tissue or background artifacts (e.g., rows 5 and 6 in Figure 1). This indicates that while VREx is designed to penalize instability, it does not guarantee perfect anatomical focus compared to the unconstrained ERM baseline.
CDD ERM Refer to caption VREx Refer to caption
CSAW Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
EMBED Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
Figure 1: Qualitative comparison of ResNet34d feature visualizations. Rows correspond to different datasets, while columns represent the training methods. Each entry displays a triplet: the original mammogram (left), the activation map for the Malignant class (center), and the activation map for the Benign class (right).
CDD ERM Refer to caption VREx Refer to caption
CSAW Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
EMBED Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
CSAW Refer to caption Refer to caption
Figure 2: Qualitative comparison of ConvNeXt-tiny feature visualizations. Rows correspond to different datasets, while columns represent the training methods. Each entry displays a triplet: the original mammogram (left), the activation map for the Malignant class (center), and the activation map for the Benign class (right).
IV Conclusion
We conducted a comprehensive evaluation of domain-invariant learning for whole-mammogram classification. Despite the theoretical appeal of methods like Invariant Risk Minimization (IRM) and Variance Risk Extrapolation (VREx), our extensive experiments on public datasets reveal their limitations in practice. IRM proved difficult to stabilize, and VREx failed to outperform the Standard ERM baseline. Our results suggest that simple aggregation of diverse training data via ERM remains the most effective strategy for this task. This study serves as a rigorous benchmark, indicating that future progress in breast cancer prediction requires novel generalization techniques capable of outperforming this strong standard baseline in out-of-domain settings.